OpenStudy (anonymous):

how do you make y+3=x+2 into standard form?

OpenStudy (anonymous):

i believe standard form is ax +by=c y-x=-1

OpenStudy (anonymous):

x-y=1

OpenStudy (anonymous):

The answer is x-y=1, you put the x first so colton is wrong.

OpenStudy (anonymous):

arrange x and y on one side and set them equal to some number. x-y=1, just as ndog33 stated.

OpenStudy (anonymous):

Sure it's not y= x-1 ?